Given the linear equation y=-3/7 x-1 , find the y-coordinates of the points (-7, ), (0, ), and (7, ). Please show all of your work. Plot those points and graph the linear equation.
Answer by rfer(16322)

y=-3/7*-7-1
y=21/7-1
y=3-1
y=2
(-7,2)
--------------
y=-3/7*0-1
y=-1
(0,-1)
-------------
y=-3/7*7-1
y=-21/7-1
y=-3-1
y=-4
(7,-4)